Grade 12Mechanics

a bullet is to be fired with speed of 20m/s to hit a target 200m away on a level ground. If g=10m/s^2, the gun should be aimed

Abhishek Talokar
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
 
Hi,

Time taken by bullet to cover 20 m horizontally = (20/200) = 0.1 s
Distance covered by bullet vertically in that time = s = 0.5 * 10 * 0.12= 0.05 m
So, the bullet should be fired 5cm above the target.
 
Regards
Arun(askIITians forum expert)